# Golden Noise for Diffusion Models: A Learning Framework

### v4

@@ 1 Introduction
Recent studies [22, 7, 5, 27, 3] observe that some selected or optimized noises are golden noises that can help the T2I diffusion models to produce images of better semantic faithfulness with text prompts, and can also improve the overall quality of the synthesized images. These methods [3, 7] incorporate extra modules like attention to reduce the truncate errors during the sampling process, showing promising results on the compositional generalization task. However, they are often not widely adopted in practice for several reasons. First, they often struggle to generally transfer to various benchmark datasets or diffusion models but only work for some specific tasks. Second, these methods often introduce significant time delays in order to optimize the noises during the reverse process. Third, they require in-depth modifications to the original pipelines when applied to different T2I diffusion models with varying architectures. Fourth, they need specific subject tokens for each prompt to calculate the loss of certain areas, which are unrealistic requirements for real users. These not only significantly complicate the original inference pipeline but also raise concerns regarding the generalization ability across various T2I diffusion models and datasets.
@@
Third, we conduct extensive experiments across various mainstream diffusion models, including StableDiffusion-xl (SDXL) [26], DreamShaper-xl-v2-turbo and Hunyuan-DiT [16], with 7 different samplers on 4 different datasets. We evaluate our model by utilizing 6 human preference metrics including Human Preference Score v2 (HPSv2) [39], PickScore [14], Aesthetic Score (AES) [34], ImageReward [41], CLIPScore [8] and Multi-dimensional Preference Score (MPS) [45]. As illustrated in Fig. 1, by leveraging the learned golden noises, not only is the overall quality and aesthetic style of the synthesized images visually enhanced, but all metrics also show significant improvements, demonstrating the effectiveness and generalization ability of our NPNet. Furthermore, the NPNet is a compact and efficient neural network that functions as a plug-and-play module, introducing only a 3% extra inference time per image compared to the standard pipeline, while requiring approximately 3% of the memory required by the standard pipeline. This efficiency underscores the practical applicability of NPNet in real-world scenarios.
@@ 3.1 Noise Prompt Dataset Collection
How to collect golden noises with desirable semantic information? Meng et al. [23] reported that adding the random noise at each timestep during the sampling process and then re-denoising, leads to a substantial improvement in the semantic faithfulness of the synthesized images. This motivated us to propose a simple and direct approach called re-denoise sampling. Instead of directly adding noise to each timestep during the reverse process, we propose to utilize $\operatorname{DDIM-Inversion}(\cdot)$ to obtain the noise from the previous step. Specifically, the joint action of DDIM-Inversion and CFG can induce the initial noise to attach semantic information. We denote the CFG scale within $\operatorname{DDIM}(\cdot)$ and $\operatorname{DDIM-Inversion}(\cdot)$ as $\omega_{l}$ and $\omega_{w}$, respectively. It is sufficient to ensure that the initial noise can be purified stably and efficiently by $\mathbf{x}^{\prime}_{t}$=DDIMInversion($\operatorname{DDIM-Inversion}(\operatorname{DDIM}(\mathbf{x}_{t}))$ with $\omega_{l}>\omega_{w}$. Utilizing this method, the synthesized image from $\mathbf{x}^{\prime}_{t}$ is more semantic information contained with higher fidelity, compared with the synthesized image from $\mathbf{x}_{t}$. We call the inverse noise $\mathbf{x}^{\prime}_{t}$ target noise, and the noise $\mathbf{x}_{t}$ source noise. The visualization results are shown in Appendix Fig. 7. The mechanism behind this method is that $\operatorname{DDIM-Inversion}(\cdot)$ injects semantic information by leveraging the CFG scale inconsistency. We present theoretical understanding of this mechanism in Theorem E.1 with the proof.
@@
Table 2, Table 3 and Appendix Table 13
Fig. 3, Appendix Fig. 10, Fig. 11 and Fig. 17
@@
Table 2, Table 3, Appendix Table 6 and Appendix Table 16
@@
Orthogonal Experiment Table 5
@@
Appendix Table 11 and Table 15
Fig. 4, Appendix Fig. 11, Fig. 13 and Fig. 17
Ablation Studies Table 4, Appendix Table 8 and Table 9
@@ 3.2 Noise Prompt Network
Model Dataset Method PickScore (↑) HPSv2 (↑) AES (↑) ImageReward (↑) CLIPScore(%) (↑) MPS(%) (↑)
Standard 21.69 28.48 6.0373 58.01 82.04 -
Inversion 21.71 28.57 6.0503 63.27 82.50 51.41
Repaint 21.77 28.63 5.9875 64.94 83.27 50.79
Pick-a-Pic NPNet (ours) 21.86 28.68 6.0540 65.01 84.08 52.14
Standard 22.31 26.72 5.5952 62.21 80.77 -
Inversion 22.37 26.91 5.6017 67.09 80.81 51.98
Repaint 22.30 26.96 5.5104 64.07 81.06 52.04
DrawBench NPNet (ours) 22.38 27.14 5.6034 70.67 81.53 53.70
Standard 22.88 29.71 5.9985 96.63 87.34 -
Inversion 22.89 29.78 5.9948 97.39 87.08 53.03
Repaint 22.91 29.78 5.9948 97.39 87.75 54.63
SDXL HPD NPNet (ours) 22.94 29.88 5.9922 98.81 88.13 56.02
Standard 22.41 32.12 6.0161 98.09 82.67 -
Inversion 22.40 32.03 6.0236 100.97 82.77 49.14
Pick-a-Pic NPNet (ours) 22.73 32.69 6.0646 106.74 89.58 52.34
Standard 22.98 30.39 5.6735 98.84 81.86 -
Inversion 22.94 30.10 5.6852 96.74 81.89 46.62
DrawBench NPNet (ours) 23.11 30.78 5.7005 108.14 82.24 53.53
Standard 23.68 30.96 6.1408 129.89 88.68 -
Inversion 23.67 31.00 6.0811 131.80 89.12 46.94
DreamShaper-xl-v2-turbo HPD NPNet (ours) 23.70 34.08 6.1283 135.98 89.42 52.49
@@ 4 Empirical Analysis
In this section, we empirically study the effectiveness, the generalization, and the efficiency of our NPNet. We conduct a lot of experiments across various datasets on various T2I diffusion models, including SDXL, DreamShaper-xl-v2-turbo, and Hunyuan-DiT. Due to space constraints, we leave implementation details and additional experiments in Appendix A and  D, respectively. We also leave the discussion and future direction in Appendix F. To assist readers in better understanding this paper, we have summarized the main tables and figures in Table. 1.
@@
We collect our NPD on Pick-a-pic dataset [14], which contains 1M prompts in its training set. We randomly choose 100k prompts as our training set. For each prompt, we randomly assign a seed value in $[0,1024]$. For testing, we use three datasets, including the first 100 prompts from the Pick-a-Pic web application, the first 100 prompts from HPD v2 test set [39], all 200 prompts from DrawBench [33] and all 553 prompts from GenEval [6]. For more details about the test data, please see in Appendix Fig. 19. We construct three training datasets collected from Pick-a-Pic, with 100k noise pairs, 80k noise pairs, and 600 noise pairs for SDXL, DreamShaper-xl-v2-turbo, and Hunyuan-DiT.
@@ 4.1 Main Results
We evaluate our NPNet in three different T2I diffusion models. The main results [^1] Method “Inversion” means re-denoise sampling. are shown in Table 2 and Table 3. In Table 2, we first evaluate our NPNet with SDXL and DreamShaper-xl-v2-turbo. The results demonstrate the impressive performance of our NPNet, almost achieving the best results across all 6 metrics and 3 datasets. Due to the space limitations, we present the experiments on GenEval with much more test prompts in Appendix D.6. We also evaluate our fine-tuned NPNet with Hunyuan-DiT, as shown in Table 3. For Hunyuan-DiT, we directly utilized the NPNet model trained on SDXL-produced NPD and fine-tune it on the Hunyuan-DiT-produced 600 noise pairs. Fine-tuned with only 600 samples, it can still achieve the highest results over the baselines on Hunyuan-DiT. This highlights the strong cross-model generalizability of our NPNet.
@@ 4.2 Analysis of Generalization and Robustness
Although we trained our NPNets exclusively with the NPDs collected from the Pick-a-Pic dataset, the experimental results presented in Table 2 and Table 3, and Fig. 3 demonstrate that our model exhibits strong cross-dataset generalization capabilities, achieving impressive results on other datasets as well. In addition to the fine-tuning experiments detailed in Table 3, we also applied the NPNet trained on NPD collected from SDXL to DreamShaper-xl-v2-turbo, evaluating the performance with our NPNet without any fine-tuning. The experiment results are shown in Appendix Table 6. These results indicate promising performance with our NPNet, underscoring the model’s capability for cross-model generalization. We also conduct experiments on noise seed range in Appendix Table 16, the results demonstrate that our NPNet exhibits strong generalization capabilities across the out-of-distribution random seed ranges.
@@
We conduct ablation studies about the architecture designs of NPNet in Table 4. The results show that both singular value prediction and residual prediction contribute to the final optimal results, while the singular value prediction component plays a more important role. We also empirically verify the effectiveness of data selection strategies in Appendix Tables 8 and 9.
@@ 5 Conclusion
In this paper, we introduce a novel concept termed the noise prompt, which aims to transform random Gaussian noise into a golden noise by incorporating a small perturbation derived from the text prompt. Building upon this, we firstly formulate a noise prompt learning framework that systematically obtains “prompted” golden noise associated with a text prompt for diffusion models, by constructing a noise prompt dataset collection pipeline that incorporates HPSv2 to filter our data and designing several backbones for our noise prompt models. Our extensive experiments demonstrate the superiority of NPNet, which is plug-and-play, straightforward, and nearly time-efficient, while delivering significant performance improvements. This model possesses considerable application potential and practical significance. We believe that the future application scope of NPNet will be broad and impactful, encompassing video, 3D content, and seamless deployment of real AIGC business products, thereby making a meaningful contribution to the community.
@@ A.1 Benchmarks
For testing, we use these four popular T2I datasets, including the first 100 prompts subset from the Pick-a-Pic web application, 100 prompts from HPD v2 test set, all 200 prompts from DrawBench and all 553 prompts from GenEval. The detailed information of the test sets is shown in Fig. 19.
@@ A.2 Evaluation Metrics
CLIPScore [8] leverages the capabilities of the CLIP model, which aligns images and text in a shared embedding space. By calculating the cosine similarity between the image and text embeddings, CLIPScore provides a mearsure of how well a synthesized image corresponds to its textual description. While CLIPScore is effective in assessing text-image alignment, it may not fully capture the nuances of human preferences, particularly in terms of aesthetic quality and detail.
@@ A.3 T2I Diffusion Models
A.4 Hyper-parameter Settings
Our method is straightforward and intuitive, and the parameter settings for the entire experiment are also very simple, with epoch 30, and batch size 64 for all experiments. We conduct experiments on three T2I diffusion models, including SDXL, DreamShaper-xl-v2-turbo, and Hunyuan-DiT, with CFG $\omega_{l}$ 5.5, 3.5, and 5.0 respectively. The inverse CFG $\omega_{w}$ is 1.0 for all three models. To collect training data, the inference steps are 10, 4, and 10 for SDXL with DDIM inverse scheduler, Dreamshaper-xl-v2-turbo with DPMSolver inverse scheduler, and Hunyuan-DiT with DDIM inverse scheduler, respectively. The human preference model we use to filter the data is the HPSv2, and the filtering threshold $k$ equals 0. Unless otherwise specified, all quantitative experiments and synthesized images in this paper are conducted and synthesize with inference steps 50, respectively. All experiments are conducted using 1$\times$ RTX 4090 GPUs, and all these noise pairs are collected with inference step 10 to construct NPDs.

### v5

@@ 1 Introduction
Recent studies [32, 11, 8, 38, 5, 2] observe that some selected or optimized noises are golden noises that can help the T2I diffusion models to produce images of better semantic faithfulness with text prompts, and can also improve the overall quality of the synthesized images. These methods [5, 11] incorporate extra modules like attention to reduce the truncate errors during the sampling process, showing promising results on the compositional generalization task. However, they are often not widely adopted in practice for several reasons. First, they often struggle to generally transfer to various benchmark datasets or diffusion models but only work for some specific tasks. Second, these methods often introduce significant time delays in order to optimize the noises during the reverse process. Third, they require in-depth modifications to the original pipelines when applied to different T2I diffusion models with varying architectures. Fourth, they need specific subject tokens for each prompt to calculate the loss of certain areas, which are unrealistic requirements for real users. These not only significantly complicate the original inference pipeline but also raise concerns regarding the generalization ability across various T2I diffusion models and datasets.
@@
Third, we conduct extensive experiments across various mainstream diffusion models, including StableDiffusion-xl (SDXL) [37], DreamShaper-xl-v2-turbo and Hunyuan-DiT [25], with 7 different samplers on 4 different datasets. We evaluate our model by utilizing 6 human preference metrics including Human Preference Score v2 (HPSv2) [53], PickScore [23], Aesthetic Score (AES) [46], ImageReward [55], CLIPScore [12] and Multi-dimensional Preference Score (MPS) [59]. As illustrated in Fig. 1, by leveraging the learned golden noises, not only is the overall quality and aesthetic style of the synthesized images visually enhanced, but all metrics also show significant improvements, demonstrating the effectiveness and generalization ability of our NPNet. Furthermore, the NPNet is a compact and efficient neural network that functions as a plug-and-play module, introducing only a 3% extra inference time per image compared to the standard pipeline, while requiring approximately 3% of the memory required by the standard pipeline. This underscores the practical applicability of NPNet in real-world scenarios.
@@ 3.1 Noise Prompt Dataset Collection
How to collect noises with desirable semantic information? Meng et al. [34] reported that adding the random noise at each timestep during the sampling process and then re-denoising, leads to a substantial improvement in the semantic faithfulness of the synthesized images. This motivated us to propose a simple and direct approach called re-denoise sampling. Instead of directly adding noise to each timestep during the reverse process, we propose to utilize $\operatorname{DDIM-Inversion}(\cdot)$ to obtain the noise from the previous step. Specifically, the joint action of DDIM-Inversion and CFG can induce the initial noise to attach semantic information. We denote the CFG scale within $\operatorname{DDIM}(\cdot)$ and $\operatorname{DDIM-Inversion}(\cdot)$ as $\omega_{l}$ and $\omega_{w}$, respectively. It is sufficient to ensure that the initial noise can be purified stably and efficiently by $\mathbf{x}^{\prime}_{t}$=DDIMInversion($\operatorname{DDIM-Inversion}(\operatorname{DDIM}(\mathbf{x}_{t}))$ with $\omega_{l}>\omega_{w}$. Utilizing this method, the synthesized image from $\mathbf{x}^{\prime}_{t}$ is more semantic information contained with higher fidelity, compared with the synthesized image from $\mathbf{x}_{t}$. We call the inverse noise $\mathbf{x}^{\prime}_{t}$ target noise, and the noise $\mathbf{x}_{t}$ source noise. The visualization results are shown in Appendix Fig. 7 and  9. The mechanism behind this method is that $\operatorname{DDIM-Inversion}(\cdot)$ injects semantic information by leveraging the CFG scale inconsistency. We present theoretical understanding of this mechanism in Theorem E.1.
@@
Table 2,  3, Appendix Table 8 and  13
Fig. 3, Appendix Fig. 12,  14 and  17
@@
Table 2,  3, Appendix Table 6,  19 and  18
@@
Orthogonal Experiment Table 5, Appendix Table 20 and Fig. 19
@@
Appendix Table 12 and  17
Fig. 4, Appendix Fig. 14,  16 and  17
Ablation Studies Table 4, Appendix Table 9 and  10
@@ 3.2 Noise Prompt Network
Model Dataset Method PickScore (↑) HPSv2 (↑) AES (↑) ImageReward (↑) CLIPScore (↑) MPS(%) (↑)
Standard 21.69 28.48 6.0373 58.01 0.8204 -
Inversion 1 21.71 28.57 6.0503 63.27 0.8250 51.41
Pick-a-Pic NPNet (ours) 21.86 28.68 6.0540 65.01 0.8408 52.14
Standard 22.31 26.72 5.5952 62.21 0.8077 -
Inversion 22.37 26.91 5.6017 67.09 0.8081 51.98
DrawBench NPNet (ours) 22.38 27.14 5.6034 70.67 0.8153 53.70
Standard 22.88 29.71 5.9985 96.63 0.8734 -
Inversion 22.89 29.78 5.9948 97.39 0.8708 53.03
SDXL HPD NPNet (ours) 22.94 29.88 5.9922 98.81 0.8813 56.02
Standard 22.41 32.12 6.0161 98.09 0.8267 -
Inversion 22.40 32.03 6.0236 100.97 0.8277 49.14
Pick-a-Pic NPNet (ours) 22.73 32.69 6.0646 106.74 0.8958 52.34
Standard 22.98 30.39 5.6735 98.84 0.8186 -
Inversion 22.94 30.10 5.6852 96.74 0.8189 46.62
DrawBench NPNet (ours) 23.11 30.78 5.7005 108.14 0.8224 53.53
Standard 23.68 30.96 6.1408 129.89 0.8868 -
Inversion 23.67 31.00 6.0811 131.80 0.8912 46.94
DreamShaper-xl-v2-turbo HPD NPNet (ours) 23.70 34.08 6.1283 135.98 0.8942 52.49
@@ 4 Empirical Analysis
In this section, we empirically study the effectiveness, the generalization, and the efficiency of our NPNet. We conduct a lot of experiments across various datasets on various T2I diffusion models, including SDXL, DreamShaper-xl-v2-turbo, and Hunyuan-DiT. Due to space constraints, we leave implementation details, related work and additional experiments in Appendix A,  C and  D, respectively. We also leave the discussion and future direction in Appendix F. To assist readers in better understanding this paper, we have summarized the main tables and figures in Table. 1.
@@
We collect our NPD on Pick-a-pic dataset [23], which contains 1M prompts in its training set. We randomly choose 100k prompts as our training set. For each prompt, we randomly assign a seed value in $[0,1024]$. For testing, we use three datasets, including the first 100 prompts from the Pick-a-Pic web application, the first 100 prompts from HPD v2 test set [53], all 200 prompts from DrawBench [45], all prompts from GenEval [10] and T2I-CompBench [20]. For more details about the test data, please see Appendix Fig. 18. We construct three training datasets collected from Pick-a-Pic, with 100k noise pairs, 80k noise pairs, and 600 noise pairs for SDXL, DreamShaper-xl-v2-turbo, and Hunyuan-DiT.
@@ 4.1 Main Results
We evaluate our NPNet in three different T2I diffusion models. The main results [^1] Method “Inversion” means re-denoise sampling . are shown in Table 2 and Table 3. In Table 2, we first evaluate our NPNet with SDXL and DreamShaper-xl-v2-turbo. The results demonstrate the impressive performance of our NPNet, almost achieving the best results across all 6 metrics and 3 datasets. Due to the space limitations, we present the experiments on GenEval and T2I-CompBench in Appendix D.7 and  8. We also evaluate our fine-tuned NPNet with Hunyuan-DiT, as shown in Table 3. For Hunyuan-DiT, we directly utilized the NPNet model trained on SDXL-produced NPD and fine-tune it on the Hunyuan-DiT-produced 600 noise pairs. Fine-tuned with only 600 samples, it can still achieve the highest results over the baselines on Hunyuan-DiT. This highlights the strong cross-model generalizability of our NPNet.
@@
We further study samples synthesized with and without NPNet to assess the distribution shift. Our empirical analysis reveals that NPNet effectively mitigates distributional discrepancies, resulting in synthesized images that exhibit statistically significant alignment with the true data manifold, shown in Appendix Fig. 13.
@@ 4.2 Analysis of Generalization and Robustness
Although we trained our NPNets exclusively with the NPDs collected from the Pick-a-Pic dataset, the experimental results presented in Table 2 and Table 3, and Fig. 3 demonstrate that our model exhibits strong cross-dataset generalization capabilities, achieving impressive results on other datasets as well. In addition to the fine-tuning experiments detailed in Table 3, we also applied the NPNet trained on NPD collected from SDXL to DreamShaper-xl-v2-turbo, evaluating the performance with our NPNet without any fine-tuning. The experiment results are shown in Appendix Table 6. Moreover, We use the same NPNet from SDXL and apply it directly on few steps T2I diffusion model, such as LCM [33], PCM [51] and SDXL-Lightning [26], which also improves the quality of the synthesized images, shown in Appendix Table 18. These results indicate promising performance with our NPNet, underscoring the model’s capability for cross-model generalization. We also conduct experiments on noise seed range in Appendix Table 19, the results demonstrate that our NPNet exhibits strong generalization capabilities across the out-of-distribution random seed ranges without harming the image diversity.
@@
More notably, we significantly improve the performance and winning rate of the previous noise optimization methods by using NPNet together with them, which further illustrates the effectiveness and scalability of our NPNet, shown in Appendix Table 20 and Fig. 19.
@@
We conduct ablation studies about the architecture designs of NPNet in Table 4. The results show that both singular value prediction and residual prediction contribute to the final optimal results, while the singular value prediction component plays a more important role. We also empirically verify the effectiveness of data selection strategies in Appendix Tables 9 and 10.
@@ 5 Conclusion
In this paper, we introduce a novel concept termed the noise prompt, which aims to transform random Gaussian noise into a golden noise by incorporating a small perturbation derived from the text prompt. Building upon this, we firstly formulate a noise prompt learning framework that systematically obtains “prompted” golden noise associated with a text prompt for diffusion models, by constructing a noise prompt dataset collection pipeline that incorporates HPSv2 to filter our data and designing several backbones for our noise prompt models. Our extensive experiments demonstrate the superiority of NPNet, which is plug-and-play, straightforward, and nearly time-efficient, while delivering significant performance improvements. We believe that the future application of NPNet will be broad and impactful, encompassing video, 3D content, and seamless deployment of AIGC products, thereby making a meaningful contribution to the community.
Acknowledgement
This work was supported by the Science and Technology Bureau of Nansha District Under Key Field Science and Technology Plan Program No. 2024ZD002. This work was supported by Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No.2023B1212010007).
@@ A.1 Benchmarks
T2I-CompBench.
T2I-CompBench is a comprehensive benchmark designed for evaluating the compositional capabilities of text-to-image (T2I) models, focusing on aspects such as object arrangement, relationships, and attributes within synthesized images. This benchmark consists of a diverse set of prompts that encompass various scenarios and contexts, facilitating a thorough assessment of how well T2I systems can interpret and visualize complex textual descriptions. By providing a structured framework for evaluation, T2I-CompBench significantly contributes to the advancement of research in T2I synthesis, offering insights into the strengths and limitations of current models in generating coherent and contextually accurate images.
For testing, we use these four popular T2I datasets, including the first 100 prompts subset from the Pick-a-Pic web application, 100 prompts from HPD v2 test set, all 200 prompts from DrawBench, all 553 prompts from GenEval and all test prompts in T2I-CompBench. The detailed information of the test sets is shown in Fig. 18.
@@ A.2 Evaluation Metrics
Figure 9: Using re-denoising sampling, we actively enhance the semantic information of the given prompt within the initial noise to enhance the semantic alignment of the generated images. This process is validated by the tight clustering of the residuals between the original noises and the inversion noises, each prompt with 100 noises, demonstrating the semantic information within the inversion noises.
@@
CLIPScore [12] leverages the capabilities of the CLIP model, which aligns images and text in a shared embedding space. By calculating the cosine similarity between the image and text embeddings, CLIPScore provides a mearsure of how well a synthesized image corresponds to its textual description. While CLIPScore is effective in assessing text-image alignment, it may not fully capture the nuances of human preferences, particularly in terms of aesthetic quality and detail222we follow https://github.com/jmhessel/clipscore.
@@ A.3 T2I Diffusion Models
A.4 Model Architecture
Our NPNet consists of two branches, one is singular value prediction, and another is residual prediction.
For the singular value prediction branch, we processes the decomposed SVD components through three parallel MLP branches: one for the orthogonal matrix U, one for V, and another for singular values s. Each branch contains two linear layers with ReLU activation. We combine them with a self-attention module to model global relationships, followed by a residual MLP that adjusts the singular values while retaining their original characteristics through skip connections. The refined components are reconstructed via matrix operations to produce the output. The detail of the self-attention is illustrated in Fig. 11.
Figure 11: The implementation details of the self-attention module in singular value prediction branch.
For the residual prediction branch, we employs a hybrid architecture combining a pretrained Swin-Tiny transformer with learnable up/down-sampling layers. The input is first upsampled to 224×224 resolution, projected to 3 channels via 1×1 convolution, then processed by the transformer’s feature extractor. Features are downsampled to the original resolution and projected back through another 1×1 convolution. An optional residual connection adds the input to the refined output.
A.5 Hyper-parameter Settings
Our method is straightforward and intuitive, and the parameter settings for the entire experiment are also very simple, with epoch 30, and batch size 64 for all experiments. During training, we surprise the obverse that randomly assigning one prompt to the whole samples in one batch can yield a better model, where we speculate that such an operation could be considered as a strong regularizer as in Chen et al. [6]. We conduct experiments on three T2I diffusion models, including SDXL, DreamShaper-xl-v2-turbo and Hunyuan-DiT, with CFG $\omega_{l}$ 5.5, 3.5, and 5.0 respectively. The inverse CFG $\omega_{w}$ is 1.0 for all three models. To collect training data, the inference steps are 10, 4, and 10 for SDXL with DDIM inverse scheduler, Dreamshaper-xl-v2-turbo with DPMSolver inverse scheduler, and Hunyuan-DiT with DDIM inverse scheduler, respectively. The human preference model we use to filter the data is the HPSv2, and the filtering threshold $k$ equals 0. Unless otherwise specified, all quantitative experiments and synthesized images in this paper are conducted and synthesize with inference steps 50, respectively. All experiments are conducted using 1$\times$ RTX 4090 GPUs, and all these noise pairs are collected with inference step 10 to construct NPDs.